OpenStudy (anonymous):

Solve the system. x=3y+9 9-x=-3y

OpenStudy (anonymous):

nonnonononon solution!

OpenStudy (anonymous):

lol what about 5x+3y=30 3x+3y=18

OpenStudy (anonymous):

x,y=(6,0)

OpenStudy (anonymous):

ok last one that's it. -10=-2y+x 6y=3x+11

OpenStudy (anonymous):

ARE YOU SURE!

OpenStudy (anonymous):

inconsistent based on the fact that the constants are not a scalar multiple of 3 therefore no solution

OpenStudy (anonymous):

yep thats it .

OpenStudy (anonymous):

did i just do like all your homework for you?;) haha least you can be proud its all right! but still you need to know whats going on!